# MultiGPrompt for Multi-Task Pre-Training and Prompting on Graphs

### 14 Dec 2023 v2

@@ E. Prompt Tuning Loss
We resort to a loss based on node/graph similarity. Consider an NC or GC task with a labeled training set $\mathcal{D}_{\text{down}}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ is either a node or a graph, and $y_{i}\in Y$ is $x_{i}$’s class label from a set of classes $Y$. Then, the prompt tuning loss is defined as $\mathcal{L}_{\text{down}}(\mathbf{\tilde{H}};\mathcal{P}_{\text{op}},\Gamma,\Delta)=$
(18) $-\sum_{(x_{i},y_{i})\in\mathcal{D}_{\text{down}}}\ln\frac{\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{\tilde{h}}_{x_{i}},\overline{\mathbf{\tilde{h}}}_{y_{i}})\right)}{\sum_{c\in Y}\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{\tilde{h}}_{x_{i}},\overline{\mathbf{\tilde{h}}}_{c})\right)},$
where $\mathbf{\tilde{h}}_{x_{i}}$ denotes the final embedding of node/graph $x_{i}$. Class $c$ prototype embedding $\overline{\mathbf{\tilde{h}}}_{c}$ is also generated based on the dual-prompt combination of composed prompt and open prompt.
Specifically, for node $v$, $\mathbf{\tilde{h}}_{v}$ is a row of $\vec{\tilde{H}}$. While for graph $G$,
(19) $\mathbf{\tilde{h}}_{G}=\textsc{Readout}(\vec{\tilde{H}}).$
F. Cross Data Scenario
Table 5. Accuracy evaluation for cross data scenario on graph classification.
@@ F. Cross Data Scenario
MultiGPrompt 60.07 $\pm$ 12.48 $\textbf{55.81}\pm 13.08$ $\textbf{59.58}\pm 13.03$ 56.17 $\pm$ 12.84
To further analyze the robustness of MultiGPrompt, we conduct additional experiments using various datasets for pre-training and downstream prompting. Specifically, we select BZR and COX2, as they possess the same node attribute dimensions, thus the pre-trained model can directly adapted to the other dataset. We respectively employ BZR and COX2 for pre-training, and perform prompting on these two datasets. The results are delineated in Table 5. We have the following observations. First, MultiGPromptconsistently surpasses other baselines even in cross-data scenario, demonstrating the robustness of the overall multi-task pre-training and prompting approach. Second, multi-task pre-training and prompting on various data do not necessarily lead to diminished performance compared to operations on identical data. This further attests to the efficacy of multi-task pre-training in acquiring effective pretext knowledge and transferring it effectively via a dual-prompt approach.

### 22 Feb 2024 v6

@@ E. Prompt Tuning Loss
We resort to a loss based on node or graph similarity. Consider an node or graph classification task with a labeled training set $\mathcal{D}_{\text{down}}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ is either a node or a graph, and $y_{i}\in Y$ is $x_{i}$’s class label from a set of classes $Y$. Then, the prompt tuning loss is defined as $\mathcal{L}_{\text{down}}(\mathbf{\tilde{H}};\mathcal{P}_{\text{op}},\Gamma,\Delta)=$
(18) $-\sum_{(x_{i},y_{i})\in\mathcal{D}_{\text{down}}}\ln\frac{\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{\tilde{h}}_{x_{i}},\overline{\mathbf{\tilde{h}}}_{y_{i}})\right)}{\sum_{y\in Y}\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{\tilde{h}}_{x_{i}},\overline{\mathbf{\tilde{h}}}_{y})\right)},$
where $\mathbf{\tilde{h}}_{x_{i}}$ denotes the final embedding of node/graph $x_{i}$, and $\overline{\mathbf{\tilde{h}}}_{y}$ is the prototype embedding of class $y$, which is the mean representation of the training instances of class $y$. $\tau$ is a temperature hyperparameter. More specifically, for a node instance $v$, $\mathbf{\tilde{h}}_{v}$ is a row of $\vec{\tilde{H}}$; for a graph instance $G$, $\mathbf{\tilde{h}}_{G}=\textsc{Readout}(\vec{\tilde{H}})$.
F. Cross-Dataset Adaptation
Table 5. Accuracy evaluation on few-shot graph classification, in same- or cross-dataset scenarios.
@@ F. Cross-Dataset Adaptation
MultiGPrompt 60.07$\pm$12.48 55.81$\pm$13.08 59.58$\pm$13.03 56.17$\pm$12.84
To further analyze the robustness of MultiGPrompt, we conduct additional experiments using different datasets for pre-training and downstream tasks, in a so-called “cross-dataset” scenario. Specifically, we select the BZR and COX2 datasets, as they possess the same node attributes, which means the model pre-trained on one dataset can be directly applied to the other dataset. We respectively employ BZR and COX2 for pre-training, and perform downstream adaptation on these two datasets.
The results for 5-shot graph classification are reported in Table 5, for same-dataset scenarios (i.e., both pretext and downstream tasks on the same dataset), as well as cross-dataset scenarios (i.e., pretext tasks on BZR and downstream tasks on COX2, or vice versa). We observe that MultiGPrompt consistently surpasses other baselines even in the cross-dataset scenarios, demonstrating the robustness of the overall multi-task pre-training and prompting strategy. Moreover, the performance of MultiGPrompt in the cross-dataset setting shows negligible decrease from the performance in the same-dataset setting. This further demonstrates the efficacy of MultiGPrompt in adapting to diverse datasets.